OpenStudy (anonymous):

What is the solution to 5y - 9 = 1? 50 40 4 2

OpenStudy (magbak):

y=2

OpenStudy (anonymous):

thanks

OpenStudy (dangerousjesse):

Solve for y: 5 y-9 = 1 Isolate terms with y to the left hand side. Add 9 to both sides: 5 y+(9-9) = 1+9 Look for two terms that sum to zero. 9-9 = 0: 5 y = 1+9 Evaluate 1+9. 1+9 = 10: 5 y = 10 Divide both sides by a constant to simplify the equation. Divide both sides of 5 y = 10 by 5: (5 y)/5 = 10/5 Any nonzero number divided by itself is one. 5/5 = 1: y = 10/5 Reduce 10/5 to lowest terms. Start by finding the GCD of 10 and 5. The gcd of 10 and 5 is 5, so 10/5 = (5×2)/(5×1) = 5/5×2 = 2:

OpenStudy (dangerousjesse):

@magbak You need to actually explain it to the student -.-

OpenStudy (magbak):

what is there to explain it is so baisic i can do it when im dying @DangerousJesse but im sorry i guess im no help and im gonna stop helping people

OpenStudy (dangerousjesse):

Well, I just explained it in full.. You need to stop taking things so seriously.

OpenStudy (magbak):

i took nothing seriously i ment like i don't correctly help people so its better for me not to complicate thing no seriousness taken im normal not sad just feel like i can do better